Circular No. 8339

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 Q4 (NEAT) C. E. Woodward, University of Minnesota; D. H. Wooden, NASA Ames

Research Center; and D. E. Harker, University of San Diego, report 8–13μm spectrophotometry of this comet on May 11.2 UT using the NASA Ames HIFOGS spectrometer at the Infrared Telescope Facility 3-m reflector: “Silicate-feature emission was observed with a flux-to-continuum ratio of 1.6, including emission from crystalline silicates at 9.3, 10.0, and 11.2 μm. The 10-μm peak flux was 1.6 × 10^{−16} W cm^{−2} μm^{−1}, and the observed N-band magnitude was −0.22 ± 0.07. A blackbody fit to the local continuum yields a color temperature of 310 ± 4 K.”

SUPERNOVAE 2004bm, 2004bp, AND 2004bq R. J. Foley, D. S. Wong, M. Ganeshalingam, A. V. Filippenko, and

R. Chornock, University of California, Berkeley, report that inspection of CCD spectra (range 330–1000 nm), obtained on May 12 UT with the Shane 3-m telescope at Lick Observatory, shows that SN 2004bm (IAUC 8335) is probably of type Ic, roughly 3–4 weeks past maximum brightness. SNe 2004bp (IAUC 8335) and 2004bq (IAUC 8336) are of type Ia, both several weeks past maximum, although the spectrum of SN 2004bp is rather noisy.

COMETS C/2003 P3, Q1, Q2, Q3, Q4, Q5, AND Q6 (SOHO) Further to IAUC 8336, K. Battams reports measurements for addi-

tional near-sun comets found on SOHO images by T. Hoffman (C/2003 P3, Q1), J. Sachs (C/2003 Q2, Q3, Q5), M. Boschat (C/2003 Q4), and R. Kracht (C/2003 Q6). All but C/2003 Q1 and Q6 (C2 images only) were observed in the C3 coronagraph; C/2003 Q2 was also observed in C2 images. The reduced astrometry by R. Kisala and orbital elements by B. G. Marsden appear on MPEC 2004-J19; C/2003 Q1 and Q6 are of the Marsden group, the other comets being Kreutz sungrazers.

Comet  2003 UT  α_{2000}  δ_{2000}
C/2003 P3 Aug. 5.763  8^{h}44.^{m}4  +15^{o}45′
C/2003 Q1  20.913  9 51.8  +13 33
C/2003 Q2  23.638  9 38.6  + 9 43
C/2003 Q3  23.846  9 46.0  +10 14
C/2003 Q4  25.113  9 55.1  +10 11
C/2003 Q5  26.279  9 47.7  + 8 54
C/2003 Q6  26.561  10 12.9  +11 42
2004 May 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green